Find the equation of the line perpendicular to 2x+3y=1 and whose distance from the
origin is 7.
1
Expert's answer
2020-10-02T15:57:41-0400

Let the meeting point of the perpendicular lines be "(a, b)"


"/OP/=\sqrt{(a^2 +b^2)} =7" ..... (i)


a2+b2=49

2x+3y=1 ;y="-(\frac{2}{3}) x+\frac{1} {3}"

Gradient ="-(\frac{2}{3})"


For perpendicular lines,the product of Gradient I and gradient II hence gradient II ="\frac{3}{2}"


"Gradient =\frac{\Delta y} {\Delta x}"


"\frac{3}{2} =\frac{b-0} {a-0}"


2b=3a .... (ii)

Substituting (ii) in (i) ;

"a^2 +(\frac{3a}{2} )^2 =49"


"4a^2+9a^2=196"


a =4 or "-"4; b=6 or "-" 6.


To get the equation, we take the point (4,6) and the gradient "\frac{3}{2}"


"\frac{3}{2} =\frac{y-6}{x-4}"


2y-12=3x -12


2y=3x


When a="-" 4 and b="-"6


"\frac{3} {2} =\frac{y-(-6)}{x-(-4)}"


3x"+" 12=2y"+"12


2y=3x